OpenStudy (anonymous):

if the national debt t years from now find relative rate of change 10 years from now in trillion of dollars N(t)=0.7+1.6e^0.03t

OpenStudy (anonymous):

n(t) = 0.6 + 1.2e^(0.01t) n'(t) = (0.01)(1.2)e^(0.01t) n'(t) = .012e^(0.01t) n'(12) = .012e^(.01(12)) n'(12) = .012e^(.12) n'(12) = 0.014 Trillions Of Dollars Per Year